OpenStudy (anonymous):

Medal Simplify (1/6)^2 - (-1/6)^2

OpenStudy (anonymous):

ok well lets put this equation together ok @harz360

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }\div2-\frac{ -1 }{ -6 }\] Correct?

OpenStudy (anonymous):

wait did that wrong srry

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

\[\left( \frac{ 1 }{ 6 } \right)\div2-\left( \frac{ -1 }{ -6 } \right)\div2\]

OpenStudy (anonymous):

there is that right?

OpenStudy (anonymous):

Where did you get the 2 from

OpenStudy (anonymous):

isnt the equation have ^2? that mean divided right?

OpenStudy (anonymous):

2 is the exponent

OpenStudy (anonymous):

oh lol srry

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

im not too good at exponents sorry ill try to get some help :)

OpenStudy (anonymous):

@dan815 @perl @BloomLocke367

OpenStudy (anonymous):

Let's look at the quantity \[(-\frac{1}{6})^2 = (-1 *\frac{1}{6})^2\] What is -1 squared?

OpenStudy (anonymous):

there you go lol ur way better than me lol

OpenStudy (anonymous):

ok how would you solve that

OpenStudy (anonymous):

What's -1 * -1?

OpenStudy (anonymous):

1

OpenStudy (shiraz14):

Use the formula a^2 - b^2 = (a-b)(a+b) ...

OpenStudy (anonymous):

Excellent. Using this, we can say \[(-\frac{1}{6})^2 = (\frac{1}{6})^2\] Can you see what it simplifies to now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we dont use the exponent?

OpenStudy (anonymous):

There are a few approaches. Shiraz gave one that is 100% valid. My approach "used" the exponent to show that we can "drop" the negative sign, since when we square a negative quantity, it's result is positive. We're then left with \[(\frac{1}{6})^2 - (\frac{1}{6})^2\] Which we can easily see goes to...

OpenStudy (anonymous):

yes now what is the next step

OpenStudy (anonymous):

We can solve it from there. We have a number (1\6)^2. We're then subtracting the same number from it. x-x = ?

OpenStudy (anonymous):

x-x = 0

OpenStudy (anonymous):

Yep. So \[(\frac{1}{6})^2 - (\frac{1}{6})^2 = ?\]

OpenStudy (kitkat16):

Squaring a number is the same as multiplying the number by itself like 6*6 so that would be 36 (1/36)-(1/6)^2 Squaring the expression is the same as multiplying the expression by itself twice 1/36((-1/6)(-1/6)) 1/36-1/36 Combine the numerators of all fractions 1/36(1-1) so subtract 1-1 and you get 0 0/36 what is an expression with zero in the numerator?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

There you go :)

OpenStudy (anonymous):

but dont the 2 minus signs become positive signs?

OpenStudy (kitkat16):

That's right 0

OpenStudy (anonymous):

The second "minus" sign is inside the squared term. We know that \[(a*b)^2 = a^2 * b^2\] so \[(-\frac{1}{6})^2 = (-1)^2*(\frac{1}{6})^2 = (\frac{1}{6})^2\]

OpenStudy (anonymous):

ohh i c so the answer is just 0?

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

thank you both i am confused on who to give the medal to lol